OpenStudy (anonymous):

Given the function f(x) = x^2+3x+2 and g(x) = x+1, determine a simplified equation for k(x)= f(g(x))

OpenStudy (kainui):

Do you know how to calculate f(2)? Just think of it the same way, but plug in g(x) instead.

OpenStudy (anonymous):

k(x)= (x^2+3x+2)(x+1) k(x)= x^3+4x^2+5x+2 <----- this the simplified answer?

OpenStudy (kainui):

Nope, what you're calculating is f(x)*g(x) but you want f(g(x)).

OpenStudy (kainui):

Pretend you want to calculate f(2) what's the first thing you do?

OpenStudy (anonymous):

imput 2 in all the x values

OpenStudy (kainui):

Perfect, so if you want to calculate f(g(x)) then just input g(x) in all the x values.

OpenStudy (anonymous):

i see... ok thanks, let me work it out

OpenStudy (kainui):

Sure, I'll tell you if you've done it right. =)

OpenStudy (anonymous):

x^2+5x+4?

OpenStudy (anonymous):

@kohai

OpenStudy (kainui):

f(x)=x^2+3x+2 so replacing all the x's with (x+1), since that's what g(x) is we get: f(g(x))=(x+1)^2+3(x+1)+2 Does that make sense?

OpenStudy (anonymous):

yah, ive done that, then i just added all like terms

OpenStudy (anonymous):

oops i made an error in my calculations

OpenStudy (anonymous):

x2+5x+6

OpenStudy (anonymous):

x^2+5x+6 is my final answer @Kainui

OpenStudy (kainui):

Yeah perfect. Good job.

OpenStudy (anonymous):

thanks you